OpenStudy (anonymous):

solve for x. x-8/2=-3

OpenStudy (anonymous):

@GeniousCreation

OpenStudy (anonymous):

x=1 x-8/2=-3 x-4=-3 1-4=-3 Hope this helps!

OpenStudy (anonymous):

Genious Creation is correct: x=1

OpenStudy (anonymous):

Yes, but you can solve x explicitly which is what they want...so x-4=-3, add 4 to both sides x = -3+4, x = 1 :))

OpenStudy (anonymous):

But note, to check it's right you let x = 1 and evaluate.